OpenStudy (anonymous):

Pleaseeee help How many mixed bags of coffee worth $90 a bag must be mixed with coffee worth $75 a bag to get 50 bags worth $87 a bag?

OpenStudy (anonymous):

let x be the no. of bags of coffee worth $90 and y the no. of bags of coffee worth $75 then, x+y=50 90x + 75y = 87*50

OpenStudy (anonymous):

Solve that two equations and find x and y... can u?